# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Acknowledgements
We thank D. Giesen, R. Hassing, S. Koot, E. Hop and T. van Klingeren for technical assistance. S.C.A. was partially supported by funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie postdoctoral fellowship No. 101106384 and the Australian Research Council Centre of Excellence for the Mathematical Modelling of Cellular Systems (MACSYS, CE230100001). J.B. acknowledges funding from the European Union’s Horizon research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No. 101106500. Y.D. acknowledges financial support from the China Scholarship Council. A.C acknowledges funding from the Research Council of Norway, project No. 39341989.
Supplementary Material: Non-reciprocal buckling makes active filaments polyfunctional
I Introduction
In this Supplementary Material, we first develop the continuum theory of non-reciprocal filaments in §II. Next we focus on the odd von Mises truss, and develop both theory (§III) and microscopic simulations (§IV) that capture the bifurcation scenario underpinning non-reciprocal buckling. Finally we describe the construction of our robotic metamaterial filaments (§V) and our experimental protocols for buckling, crawling, digging and walking (§VI).
II Continuum theory
@@
where the last step has used the inextensibility condition, Eq. (S4). Similarly, for the change in normal vector and curvature:
@@
$\beta U$ $=-\Lambda\kappa+A\left(\partial_{s}^{2}\kappa+\frac{1}{2}\kappa\left(\kappa^{2}-\kappa_{0}^{2}\right)\right)+\zeta\partial^{3}_{s}\kappa+\eta\partial_{s}^{3}(W\kappa-\partial_{s}U),$ (18)
@@
First we consider the case of an infinite line parameterised by ${\bf x}=(x(s,t),h(s,t))$ where $h(s,t)\ll 1$. At zeroth order in height deformations the filament basis vectors are given by ${\bf t}={\bf e}_{x}$ and ${\bf n}={\bf e}_{y}$. The curvature is given by $\kappa=-h_{xx}$ and $U=\dot{h}+U_{0}$. For simplicity we will assume that the spontaneous curvature is zero ($\kappa_{0}=0$). The continuum dynamics Eq. (19) are solved exactly for $U=W=h=0$ with $\Lambda=\text{Const.}=\Lambda_{0}+\delta\Lambda$. We will perturb around this base solution.
@@
where $\lambda$ is the growth rate of the spatial Fourier mode with wavenumber $q$. Equation (23) appears as Eq. (2) of the Main Text.
III Analysis of the odd von Mises truss
Derivation of the equations of motion via the principle of virtual work
The odd active torque for each of the vertices is given by
$\tau_{\text{L}}$ $=k^{\text{o}}(\theta_{\text{LC}}-\pi)\text{,}$ (24)
$\tau_{\text{R}}$ $=-k^{\text{o}}(\theta_{\text{RC}}-\pi)\text{,}$ (27)
$E_{s}=$ $\frac{1}{2}k\left(l_{\text{0}}-\sqrt{(-\text{d}l+l-w_{\text{L}}-w_{\text{R}})^{2}+(w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}}))^{2}}\right)^{2}$
$\mathcal{R}=\frac{\Gamma}{2}\left[\dot{\theta}_{\text{L}}^{2}+\dot{\theta}_{\text{LC}}^{2}+\dot{\theta}_{\text{RC}}^{2}+\dot{\theta}_{\text{R}}^{2}\right]+k^{\text{o}}\left[(\theta_{\text{LC}}-\pi)\dot{\theta}_{\text{L}}+(\theta_{\text{RC}}-\pi-\theta_{\text{L}})\dot{\theta}_{\text{LC}}+(\theta_{\text{R}}-\theta_{\text{LC}}+\pi)\dot{\theta}_{\text{RC}}-(\theta_{\text{RC}}-\pi)\dot{\theta}_{\text{R}}\right]\text{.}$ (30)
Here $\Gamma$ is the hinge dissipation and $k^{\text{o}}$ the microscopic non-reciprocity. Terms like $\frac{\Gamma}{2}\dot{\theta{L}}^{2}$ in Eq. (30) are the viscous power dissipated by hinge $\theta_{L}$. Terms like $k^{\mathrm{o}}(\theta_{\text{LC}}-\pi)\dot{\theta}_{\text{L}}$ are the power generated by the odd torques acting on vertex $\theta_{\text{L}}$. Given a Lagrangian $L(\theta,\dot{\theta})$ that encodes conservative dynamics, and a dissipation function $\mathcal{R}(\theta,\dot{\theta})$, the Rayleighan formalism gives equation of motion via
In the truss, the Lagrangian $L=E_{s}+E_{b}$ is the sum of bending and stretching energies Eqs. (28), (29), neglecting inertia. The dissipation $\mathcal{R}$ is given by Eq. (30).
$\theta_{\text{LC}}=\pi-\theta_{\text{L}}-\text{arctan}\left(-\text{d}l+l-w_{\text{L}}-w_{\text{R}},w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}})\right)\text{,}$ (32)
$\theta_{\text{RC}}=\pi-\theta_{\text{R}}+\text{arctan}\left(-\text{d}l+l-w_{\text{L}}-w_{\text{R}},w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}})\right)\text{.}$ (33)
We also non-dimensionalize energies by $B$ and lengths by $l=3l_{0}$—note that this is a different lengthscale to the discrete simulations by a factor of 3. Time will be renormalized by $\Gamma/B$. Thus our dimensionless parameters are given by
$k^{\text{o}}\leftarrow\frac{k^{\text{o}}}{B}\text{,}$ (38)
$k\leftarrow\frac{kl^{2}}{B}\text{,}$ (39)
$\text{d}l\leftarrow\frac{\text{d}l}{l}\text{.}$ (40)
Expanding to $4^{\text{th}}$ order in $\epsilon$ gives the following energy:
$+\frac{1}{576}\epsilon^{4}\Big(144\text{d}l^{2}k-2592\text{d}l\theta_{A}^{2}-24\text{d}l\theta_{A}^{2}k-96\text{d}l\theta_{S}^{2}k-288\text{d}lkw_{\text{S}}-432\theta_{A}^{4}+1728\theta_{A}^{2}\theta_{S}^{2}+9\theta_{A}^{4}k$
$\qquad+24\theta_{A}^{2}\theta_{S}^{2}k+16\theta_{S}^{4}k+144kw_{\text{A}}^{2}-96\theta_{A}\theta_{S}kw_{\text{A}}+432kw_{\text{S}}^{2}-72\theta_{A}^{2}kw_{\text{S}}+96\theta_{S}^{2}kw_{\text{S}}-5184\theta_{A}\theta_{S}w_{\text{A}}$
$\qquad+7776\theta_{A}^{2}w_{\text{S}}\Big)\text{,}$ (41)
and the dissipation is given by
$\mathcal{R}=$ $\frac{1}{2}\epsilon^{2}\left(10k^{\text{o}}\theta_{\text{S}}\dot{\theta}_{\text{A}}-10k^{\text{o}}\theta_{\text{A}}\dot{\theta}_{\text{S}}+5\dot{\theta}_{\text{A}}^{2}+4\dot{\theta}_{\text{S}}^{2}\right)+\frac{1}{4}\epsilon^{4}\Big(-72\text{d}lk^{\text{o}}\theta_{\text{S}}\dot{\theta}_{\text{A}}+72\text{d}lk^{\text{o}}\theta_{\text{A}}\dot{\theta}_{\text{S}}-54\text{d}l\dot{\theta}_{\text{A}}^{2}-12k^{\text{o}}\theta_{\text{A}}^{2}\theta_{\text{S}}\dot{\theta}_{\text{A}}$
$\qquad+16k^{\text{o}}\theta_{\text{S}}^{3}\dot{\theta}_{\text{A}}+4k^{\text{o}}\theta_{\text{A}}^{3}\dot{\theta}_{\text{S}}+16k^{\text{o}}\theta_{\text{A}}\theta_{\text{S}}^{2}\dot{\theta}_{\text{S}}-48k^{\text{o}}\theta_{\text{S}}^{2}\dot{w}_{\text{A}}+72k^{\text{o}}\theta_{\text{A}}\theta_{\text{S}}\dot{w}_{\text{S}}+72k^{\text{o}}\theta_{\text{S}}w_{\text{S}}\dot{\theta}_{\text{A}}-72k^{\text{o}}\theta_{\text{A}}w_{\text{S}}\dot{\theta}_{\text{S}}$
$\qquad+24\theta_{\text{A}}\theta_{\text{S}}\dot{\theta}_{\text{A}}\dot{\theta}_{\text{S}}+12\theta_{\text{S}}^{2}\dot{\theta}_{\text{A}}^{2}-9\theta_{\text{A}}^{2}\dot{\theta}_{\text{A}}^{2}-36\theta_{\text{S}}\dot{\theta}_{\text{A}}\dot{w}_{\text{A}}-36w_{\text{A}}\dot{\theta}_{\text{A}}\dot{\theta}_{\text{S}}+54\theta_{\text{A}}\dot{\theta}_{\text{A}}\dot{w}_{\text{S}}+54w_{\text{S}}\dot{\theta}_{\text{A}}^{2}\Big)\text{.}$ (42)
$\frac{\partial}{\partial w_{\text{{S,A}}}}(E_{b}+E_{s})=0,$ (43)
which gives
$w_{\text{A}}=\frac{(k+54)\theta_{\text{A}}\theta_{\text{S}}}{3k},$ (44)
We then set $\dot{w}_{\text{S,A}}=0$ in the dissipation functional Eq. (42) and calculate the dynamical equations for only the two angular degrees of freedom. These equations are found by application of the variational principle Eq. (31) as
which to third order in $\epsilon$ gives
Equations (47) are of the general cubic form given in Eq. (3) of the Main Text, namely
@@ III Analysis of the odd von Mises truss
As described in the Main Text, we note that the $\mathbb{Z}_{2}$ symmetry $(\theta_{S},\theta_{A})\rightarrow(-\theta_{S},-\theta_{A})$ forces nonlinear terms of cubic order only in Eq. (48). The symmetry under simultaneous reflection and a reversal in the sense of non-reciprocity, $(\theta_{S},\theta_{A},k^{\mathrm{o}})\rightarrow(\theta_{S},-\theta_{A},-k^{\mathrm{o}})$ constrains $k^{\mathrm{o}}$ to appear only in the off-diagonal elements of $J$ in Eq. (49), and enforces the alternating pattern of terms proportional to $k^{\mathrm{o}}$ in ${\bf A}_{i}$. More formally, we split these terms into ${\bf A}_{i}={\bf A}_{i}^{\mathrm{passive}}(k)+{\bf A}_{i}^{\mathrm{odd}}(k^{\mathrm{o}},k)$. We require the symmetry ${\bf A}^{\mathrm{odd}}_{i}=(-1)^{i}\sigma_{z}{\bf A}^{\mathrm{odd}}_{i}$, with $\sigma_{z}$ a Pauli matrix. We now analyze the bifurcation structure of the dynamics Eq. (48).
Bifurcation structure of the odd von Mises truss
Figure S1: Phase portraits of the odd von Mises truss. Phase portraits of the dynamics of the odd von Mises truss Eq. (48) in $\theta_{\text{S}}$ (horizontal) and $\theta_{\text{A}}$ (vertical) space, for different values of compression, $\text{d}l$, and non-reciprocity, $k^{\text{o}}$. Blue dots indicate fixed points. For a given $\text{d}l$, as $k^{\mathrm{o}}$ increases fixed points representing buckled states coalesce to form a limit cycle describing self-snapping.
Figure S1 shows a selection of phase portraits for the truss dynamics Eq. (48) as the non-reciprocity $k^{\mathrm{o}}$ and compression $\text{d}l$ are varied; an annotated subset of these phase diagrams are shown in Fig. 3(e) of the Main Text. In the limit of no active force, $k^{\text{o}}=0$, we find classic symmetrically buckled solutions
$\theta_{\text{S}}^{\star}=\pm\frac{\sqrt{3}\sqrt{\text{d}lk-18}}{\sqrt{k}},\quad\theta_{\text{A}}^{\star}=0\text{,}$ (50)
where are plotted in Fig. S2 (c.f. Fig. S1, top row) . As $k^{\mathrm{o}}$ is increased, these buckled fixed points merge in a SNIC bifurcation, and a limit cycle is born (Fig. S1, bottom row). Example numerical trajectories in $\theta_{\text{{S,A}}}$ as $k^{\mathrm{o}}$ increases are shown in Fig. S3.
Figure S2: Passive buckling of the odd von Mises truss. When non-reciprocity $k^{\mathrm{o}}=0$, the dynamics of the odd von Mises truss Eq. (48) recovers the pitchfork bifurcation of classic buckling. We show the evolution of fixed points for the symmetric angle $\theta_{\text{S}}^{\star}$ as a function of compression $\text{d}l$ for $k^{\text{o}}=0$, $k=1000$.
Figure S3: Self-Snapping trajectories. Above the threshold $\text{d}l_{c}$ for buckling Eq. (54), as $k^{\mathrm{o}}$ is increased, the truss initially polarizes ($\theta_{A}$ increases) and then begins to self-snap. Trajectories shown in $\theta_{\text{S,A}}$ come from numerical integration of Eq. (48) at $\text{d}l=0.05$, $k=1000$ and $k^{\text{o}}=0,0.1,1$ from left to right.
To locate the critical non-reciprocity $k_{c}^{o}$ and compression $\text{d}l_{c}$ for the CEP shown in Fig. 3 of the Main Text, we demand a double zero in the eigenvalues of the Jacobian $J$ in Eq. (49). We enforce this condition by solving for $\text{Tr}({J})=\text{Det}({J})=0$, which gives
$k^{\text{o}}_{\text{c}}=\pm\frac{k+162}{\sqrt{20k^{2}-2160k+25515}},$ (51)
$\text{d}l_{\text{c}}=\frac{45}{2k-81}\text{.}$ (52)
$\text{d}l_{\text{c}}=\frac{45}{2k}\text{,}$ (54)
which is the location of the CEP given in the Main Text. Substituting these critical values back into the Jacobian Eq. (49), perturbing as $k^{\text{o}}=k^{\text{o}}_{\text{c}}+\delta k^{\text{o}}$, $\text{d}l=\text{d}l_{\text{c}}+\delta l/k$ and taking the limit $k\to\infty$ we find the normal form
which is given as Eq. 4 of the Main Text. In order to find the theoretical snapping frequencies shown in Fig. 3(b) of the Main Text we look for the imaginary part of eigenvalues, $\lambda_{J}$ of $J$ in the vicinity of the CEP:
Figure S4: Structure of the critical exceptional point. (a) The linearized dynamics about the CEP in Eq. (48). Colours denote the structure of the fixed point at the origin, stable aster (red), saddle (green), unstable aster (organge), stable spiral (blue) and unstable spiral (purple), calculated according to the eigenvalues of $J$ at the origin. (b) Structure of the critical exceptional point in $\mathbb{Z}_{2}$ symmetric Bogdanov-Takens bifurcation. Colors are calculated as in (a). (c) Zoom in of the full system critical point, showing the region where the $\mathbb{Z}_{2}$ Bogdanov-Takens bifurcation occurs. Each line corresponds to a different type of bifurcation (for simplicity saddle connection bifurcations are omitted). Black: pitchfork bifurcation, blue: Hopf at origin, green: global saddle-node of two cycles, purple: symmetric Hopf bifurcations not at origin, red: saddle node bifurcation. (d) Phase plots showing the $(\theta_{\mathrm{S}},\theta_{\mathrm{A}})$ plane for the two paths shown in (c), that is varying $k^{\mathrm{o}}$ while leaving $\mathrm{d}l$ fixed. The colored lines between phase portraits correspond to the lines denoting bifurcations in panel (c).
$\mathbb{Z}_{2}$ symmetric Bogdanov-Takens bifurcation
The CEP in the odd von Mises truss is a single point in parameter space, which occurs as part of a codimension-2 bifurcation scenario called a $\mathbb{Z}_{2}$ Symmetric Bogdanov-Takens bifurcation [37]. This bifurcation has normal form
$y_{1}^{\prime}=y_{2}$ (57)
with Jacobian given by
@@
In Fig. S4(a, b) we compare the Jacobian of the truss Eq. (49) to that of the normal form Eq. (59), showing that they are related by a smooth deformation. For each case we compute the eigenvalues of the Jacobian for the fixed point at the origin and categorize into the following standard cases. Stable aster [two negative real eigenvalues, red in Fig. S4(a,b)], saddle [one negative and one positive real eigenvalue, green in Fig. S4(a,b)], unstable aster [two positive real eigenvalues, orange in Fig. S4(a,b)], unstable spiral [two complex eigenvalues with positive real part, purple in Fig. S4(a,b)], and stable spiral [two complex eigenvalues with negative real part, blue in Fig. S4(a,b)].
@@
IV Discrete simulations
In this section, we perform additional discrete simulations of non-reciprocal beams which allow us to assess the robustness of assumptions made in our theoretical modelling. Our truss theory (§III) assumes small strain, and neglects inertia. Here we numerically validate these assumptions.
@@ IV Discrete simulations
where $\epsilon$ is the Levi-Civita symbol. The first terms in Eq. (61) are the forces due to the angular tensions, while the second terms are the Hookean resistance to stretching. The functions $\tau_{\alpha}(\underline{\mathbf{x}})$ determine the vertex bending properties. For our non-reciprocal torsional elements, we model the vertices via
@@
For convenience, we list all microscopic and continuum parameters, alongside their dimension, in Table 1. On dimensional grounds, and comparing to the linearized theory Eq. (22), we can making the following correspondence between the microscopic and continuum parameters:
@@
Name Units
System Size $L$ $L$
Microscopic
mass $m$ $M$
Lattice Spacing $l_{0}$ $L$
Bending Rigidity $B$ $E$
Longitudinal spring stiffness $k$ $EL^{-2}$
Microscopic Non-reciprocity $k^{\text{o}}$ $E$
Hinge dissipation $\Gamma$ ETET
Continuum
Line density $\rho$ $EL^{-1}$
Bending Modulus AA ELEL
Continuum Non-reciprocity $\zeta$ $EL^{2}$
Bending Viscosity $\eta$ ELTELT
Table 1: Names and units of material parameters.
@@
Figure S6: Effect of finite mass and compressive strain on the self-snapping instability. Our theory Eq. (48) is developed in the limit that $\tilde{m}\rightarrow 0$ and of small compressive strain. However, our numerics show that our theory quantitatively captures the instability at $O(1)$ values of both parameters. In panel (a), $\tilde{m}=0.01$. In panel (b), strain is $0.1$. Throughout these simulations $\tilde{\beta}=0$, $\tilde{k}=1000$, with boundary compression implemented via stiff longitudinal springs with $\tilde{k}=5000$.
We non-dimensionalise Eq. (61) with a viscous timescale, and use the length of a single linkage as a characteristic length scale:
@@
These length and time scales are identical to those used in our microscopic theory of the von Mises truss §III, except that we non-dimensionalize by a single lattice spacing $l_{0}$ rather than the entire system size $l=Nl_{0}$. The resulting dynamics is described by five dimensionless groups that mirror those used in our microscopic theory §III,

### v2

@@ Acknowledgements
All the codes and data supporting this study are available on the public repository at 10.5281/zenodo.17513595. S.C.A.-I. and Y.D contributed equally to this work. S.C.A.-I., Y.D., J.V., A.S., A.C., and C.C. and J.B. designed research; S.C.A.-I., Y.D., J.V., R.G.M., A.C., C.C., and J.B. performed research; S.C.A.-I., Y.D., J.V., and J.B. analyzed data; R.G.M., A.S., and A.C. feedback on paper; and S.C.A.-I., C.C., and Y.D. and J.B. wrote the paper. We thank D. Giesen, R. Hassing, S. Koot, E. Hop, and T. van Klingeren for technical assistance. S.C.A.-I. was partially supported by funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie postdoctoral fellowship No. 101106384 and the Australian Research Council Centre of Excellence for the Mathematical Modelling of Cellular Systems (MACSYS, CE230100001). J.B. acknowledges funding from the European Union’s Horizon research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No. 101106500. Y.D. acknowledges financial support from the China Scholarship Council. A.C. acknowledges funding from the Research Council of Norway, project No. 39341989 and from the European Research Council under grant agreement ERC-CoG 101169717. C.C. acknowledges funding from the Netherlands Organization for Scientific Research under grant agreement VI.Vidi.213.131 and from the European Research Council under grant agreement ERC-CoG 101170693.
Supplementary Information: Nonreciprocal buckling makes active filaments polyfunctional
V Introduction
In this Supplementary Information, we first develop the continuum theory of non-reciprocal filaments in §VI. Next we focus on the odd von Mises truss. We first provide more details of the $\mathbb{Z}_{2}$ Bogdanov-Takens bifurcation, which contains the CEP in its linearized dynamics (§VII). Next we develop microscopic simulations (§VIII) that complement the theory described in the Main Text and capture the bifurcation scenario underpinning non-reciprocal buckling. Finally, we show that a critical exceptional point also underpins the buckling of an entirely different class of active beam composed of follower forces (§IX).
VI Continuum theory
Name Symbol Units
System Size $l$ $L$
Microscopic
Mass $m$ $M$
Lattice Spacing $l_{0}$ $L$
Bending Rigidity $B$ $E$
Longitudinal Spring Stiffness $k$ $EL^{-2}$
Microscopic Non-reciprocity $k^{\text{o}}$ $E$
Hinge Dissipation $\Gamma$ ETET
Continuum
Line Density $\rho$ $EL^{-1}$
Bending Modulus AA ELEL
Continuum Non-reciprocity $\zeta$ $EL^{2}$
Bending Viscosity $\eta$ ELTELT
Table 1: Names, symbols, and units of material parameters.
@@ VI Continuum theory
where the last step has used the inextensibility condition, (4). Similarly, for the change in normal vector and curvature:
@@
$\beta U$ $=-\Lambda\kappa+A\left(\partial_{s}^{2}\kappa+\frac{1}{2}\kappa\left(\kappa^{2}-\kappa_{0}^{2}\right)\right)$
$\quad+\zeta\partial^{3}_{s}\kappa+\eta\partial_{s}^{3}(W\kappa-\partial_{s}U),$ (18)
@@
First we consider the case of an infinite line parameterised by ${\bf x}=(x(s,t),h(s,t))$ where $h(s,t)\ll 1$. At zeroth order in height deformations the filament basis vectors are given by ${\bf t}={\bf e}_{x}$ and ${\bf n}={\bf e}_{y}$. The curvature is given by $\kappa=-h_{xx}$ and $U=\dot{h}+U_{0}$. For simplicity we will assume that the spontaneous curvature is zero ($\kappa_{0}=0$). The continuum dynamics (19) are solved exactly for $U=W=h=0$ with $\Lambda=\text{Const.}=\Lambda_{0}+\delta\Lambda$. We will perturb around this base solution.
@@
where $\lambda$ is the growth rate of the spatial Fourier mode with wavenumber $q$. (23) appears as Eq. (3) of the Main Text.
VII The $\mathbb{Z}_{2}$ symmetric Bogdanov-Takens bifurcation
Figure S1: Structure of the critical exceptional point. (a) The linearized dynamics about the CEP in (27). Colours denote the structure of the fixed point at the origin, stable aster (red), saddle (green), unstable aster (organge), stable spiral (blue) and unstable spiral (purple), calculated according to the eigenvalues of $J$ at the origin. (b) Structure of the critical exceptional point in $\mathbb{Z}_{2}$ symmetric Bogdanov-Takens bifurcation. Colors are calculated as in (a). (c) Zoom in of the full system critical point, showing the region where the $\mathbb{Z}_{2}$ Bogdanov-Takens bifurcation occurs. Each line corresponds to a different type of bifurcation (for simplicity saddle connection bifurcations are omitted). Black: pitchfork bifurcation, blue: Hopf at origin, green: global saddle-node of two cycles, purple: symmetric Hopf bifurcations not at origin, red: saddle node bifurcation. (d) Phase plots showing the $(\theta_{\mathrm{S}},\theta_{\mathrm{A}})$ plane for the two paths shown in (c), that is varying $k^{\mathrm{o}}$ while leaving $\mathrm{d}l$ fixed. The colored lines between phase portraits correspond to the lines denoting bifurcations in panel (c).
The CEP in the odd von Mises truss is a single point in parameter space, which occurs as part of a codimension-2 bifurcation scenario called a $\mathbb{Z}_{2}$ Symmetric Bogdanov-Takens bifurcation [23]. This bifurcation has a normal form
$y_{1}^{\prime}=y_{2}$ (24)
with Jacobian given by
By contrast, in the Main Text we derive the dynamics of our truss dynamics as:
@@ VII The $\mathbb{Z}_{2}$ symmetric Bogdanov-Takens bifurcation
In Fig. S1(a, b) we compare the Jacobian of the truss (28) to that of the normal form (26), showing that they are related by a smooth deformation. For each case we compute the eigenvalues of the Jacobian for the fixed point at the origin and categorize into the following standard cases. Stable aster [two negative real eigenvalues, red in Fig. S1(a, b)], saddle [one negative and one positive real eigenvalue, green in Fig. S1(a, b)], unstable aster [two positive real eigenvalues, orange in Fig. S1(a, b)], unstable spiral [two complex eigenvalues with positive real part, purple in Fig. S1(a, b)], and stable spiral [two complex eigenvalues with negative real part, blue in Fig. S1(a, b)].
@@
VIII Discrete simulations
In this section, we perform additional discrete simulations of non-reciprocal beams which allow us to assess the robustness of assumptions made in our theoretical modelling. Our truss theory in the Main Text assumes small strain, and neglects inertia. Here we numerically validate these assumptions.
@@ VIII Discrete simulations
where $\epsilon$ is the Levi-Civita symbol. The first terms in (30) are the forces due to the angular tensions, while the second terms are the Hookean resistance to stretching. The functions $\tau_{\alpha}(\underline{\mathbf{x}})$ determine the vertex bending properties. For our non-reciprocal torsional elements, we model the vertices via
@@
For convenience, we list all microscopic and continuum parameters, alongside their dimension, in Table 1. On dimensional grounds, and comparing to the linearized theory (22), we can make the following correspondence between the microscopic and continuum parameters:
@@
Figure S3: Effect of finite mass and compressive strain on the self-snapping instability. Our theory (27) is developed in the limit that $\tilde{m}\rightarrow 0$ and of small compressive strain. However, our numerics show that our theory quantitatively captures the instability at $O(1)$ values of both parameters. In panel (a), $\tilde{m}=0.01$. In panel (b), strain is $0.1$. Throughout these simulations $\tilde{\beta}=0$, $\tilde{k}=1000$, with boundary compression implemented via stiff longitudinal springs with $\tilde{k}=5000$.
We non-dimensionalise (30) with a viscous timescale, and use the length of a single linkage as a characteristic length scale:
@@
These length and time scales are identical to those used in our microscopic theory of the von Mises truss, except that we non-dimensionalize by a single lattice spacing $l_{0}$ rather than the entire system size $l=Nl_{0}$. The resulting dynamics is described by five dimensionless groups that mirror those used in our microscopic theory,